Go to the St. Louis Federal Reserve FRED database, and find the June 2013 data available...

  1. Go to the St. Louis Federal Reserve FRED database, and find the June 2013 data available on Currency (CURRNS), Total Checkable Deposits (TCDNS), Total Reserves (RESBALNS), and Required Reserves (RESBALREQ).
    1. Calculate the value of the currency deposit ratio (c).
    2. Use RESBALNS and RESBALREQ to calculate the amount of excess reserves, and then calculate the value of the excess reserve ratio (e). Be sure the units of total and required reserves are the same when you do the calculations.
    3. Assuming a required reserve ratio (rr) of 11%, calculate the value of the money multiplier (m).
Answer #1

1.

COMPUTATIONAL TABLE

Variables

CURRNS

TCDNS

RESBALNS

RESBALREQ

Values in US$ billions (as on 1st July, 2013)*

1130.3

1409.8

2094.225

63.296

*NOTE: Data was not available for July 31st, 2013

a. Currency deposit ratio (c) = CURRNS/TCDNS = (1130.3/1409.8) = 0.8017

b. Excess Reserves (ER) = RESBALNS – RESBALREQ = 2030.929

    Excess Reserves Ratio (e) = ER/TCDNS = 1.4405

c. Money Multiplier (m) = (1 + c)/(c + rr) = 1.9761